# White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs

### v3

@@ Abstract
Language agency is an important aspect of evaluating social biases in texts. While several studies approached agency-related bias in human-written language, very limited research has investigated such biases in Large Language Model (LLM)-generated content. In addition, previous research often relies on string-matching techniques to identify agentic and communal words within texts, which fall short of accurately classifying language agency. We introduce the novel Language Agency Bias Evaluation (LABE) benchmark, which comprehensively evaluates biases in LLMs by analyzing agency levels attributed to different demographic groups in model generations. LABE leverages 5,400 template-based prompts, an accurate agency classifier, and corresponding bias metrics to test for gender, racial, and intersectional language agency biases in LLMs on 3 text generation tasks: biographies, professor reviews, and reference letters. To build better and more accurate automated agency classifiers, we also contribute and release the Language Agency Classification (LAC) dataset, consisting of 3,724 agentic and communal sentences. Using LABE, we unveil previously under-explored language agency social biases in 3 recent LLMs: ChatGPT, Llama3, and Mistral. We observe that: (1) For the same text category, LLM generations tend to demonstrate higher levels of gender bias than human-written texts; (2) On most generation tasks, models demonstrate remarkably higher levels of intersectional bias than the other bias aspects. Those who are at the intersection of gender and racial minority groups—such as Black females—are consistently described by texts with lower levels of agency, aligning with real-world social inequalities; (3) Among the 3 LLMs investigated, Llama3 demonstrates greatest overall bias in language agency; (4) Not only does prompt-based mitigation fail to resolve language agency bias in LLMs, but it frequently leads to the exacerbation of biases in generated texts.
@@ 1 Introduction
Social biases manifest through the varying levels of agency in texts describing different demographic groups Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023). For instance, bias exists in texts portraying demographic minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020). While a body of works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022; Wan et al., 2023) have studied agency level in texts, these previous works suffer from several remarkable drawbacks:
@@
2. Existing methods to measure language agency struggle with achieving accuracy and reliability. Prior works often utilized string matching for words in agentic and communal lexicons to measure agency. However, string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively (as shown in Appendix C, Table 11). A qualitative example is provided in Figure 1: while differences in language agency are clearly observable in the texts, string matching yields 00 agentic and 00 communal words for both texts; sentiment classifier labels both sentences as “positive”.  Wan et al. (2023) utilized a model-based method for measuring agency, but their model only achieves $66.49\%$ classification accuracy (Appendix C, Table 11).
@@ 2.1 Biases in Human-Written and LLM-generated Texts
The presence of gender, racial, and intersectional bias in human society has significantly impacted human language Blodgett et al. (2020); Doughman et al. (2021) and generative LLMs, which utilize extensive texts for training. In this work, we investigate biases in 3 different categories of texts: biographies, professor reviews, and reference letters.
Bias in BiographiesWagner et al. (2016); 10.1145/3485447.3512134, and  Park et al. (2021) studied gender biases in Wikipedia biographies.  Park et al. (2021) analyzed biases in power, agency, and sentiment words in biography pages;  Wagner et al. (2016) revealed negative linguistic biases in womens’ pages.  10.1145/3485447.3512134 and  Adams et al. (2019) studied racial biases in editorial traits such as length and academic rank.  10.1145/3485447.3512134; Adams et al. (2019) and  Lemieux et al. (2023) stressed the importance of studying intersectional gender and racial biases in Wikipedia. Along similar lines, Otterbacher (2015) found biases towards Black female actresses in IMDB biographies. Bias in Professor Reviews   Prior works (Roper, 2019; Macnell et al., 2014) have revealed gender biases in student ratings for professors—instructors with female perceived gender received lower ratings than males. Schmidt visualized the gendered language in RateMyProfessor reviews by string matching for gender-indicative words.  Reid (2010) showed that professors from racial minority groups received more negative RateMyProfessor evaluations.  Chávez and Mitchell (2020) further revealed intersectional gender and racial biases towards female professors from racial minority groups in professor reviews.
@@ 3 The Language Agency Bias Evaluation (LABE) Framework
In this paper, we propose the Language Agency Bias Evaluation (LABE) benchmark for comprehensively assessing language agency biases in LLMs across race, gender, and intersectional identities. LABE prompts an LLM to generate descriptive texts for multiple demographic groups, and assesses biases by inspecting the variability in language agency between generations for different groups.
Split # Entries # Agentic # Communal
Train $2,979$ $1,523$ $1,456$
Valid $372$ $192$ $180$
Test $373$ $199$ $174$
Table 1: LAC Dataset Statistics.
Text Type Template # Entries
Table 2: Templates for LLM Generations in LABE.
@@ 3.1 Generative Discriptive Texts for Demographic Groups with LLMs
Wan et al. (2023) proposed the Context-Less Generation (CLG) setting, in which they adopt templates and descriptors to prompt for a variety of LLM-generated reference letters for different genders. Inspired by CLG, we extend the setting to 3 different text generation tasks: biography, professor review, and reference letter generation. We combine descriptors with demographic information—such as race, gender, or intersectional identities—and template-based prompts to query for LLMs’ generation. Each prompt must contain race and gender descriptors. For the name descriptor, we prompt ChatGPT to generate 5 popular names for each gender and race intersectional group. Descriptors for additional details like occupation and department are included to improve prompt variability. The final LABE benchmark tests LLMs on 2,400 templated-based prompts for biography generation, 600 for professor review, and 2,400 for reference letters. Note that entry numbers differ due to the difference in descriptors used (departments for professor review, whereas occupations for the other 2). We provide examples of each descriptor below, with full details included in Appendix A:
  1. 1.
    Race: “Black”, “White”, “Hispanic”, “Asian”
  2. 2.
    Gender: “Male”, “Female”
  3. 3.
    Name:
    1. (a)
      “Asian” + “Female”: “Mei”, “Aiko”, “Linh”, “Priya”, “Ji-Yoon”
    2. (b)
  4. 4.
    Occupation: “student”, “entrepreneur”,“actor”, “artist”, “chef”, …
  5. 5.
    Department: “Communication department”, “Fine Arts department”, “Chemistry department”, …
@@ 3.2 Evaluating Language Agency: The Language Agency Classification (LAC) Dataset
For building accurate automated evaluation tools for language agency, we propose the Language Agency Classification (LAC) dataset, a corpus with 3,724 agentic and communal sentences with corresponding labels. The dataset construction process incorporates an efficient automated generation pipeline, and careful verification by human annotators who are native speakers of English.
@@ 3.2.1 Dataset Collection
    The data construction process should be scalable. Since we are constructing a classification dataset, we need to ensure enough entries for agentic language and communal language to train a useful classifier.
@@
    The data construction process should be consistent. Quality of texts for all entries should remain consistent, and should not differ from part to part.
@@
    The data construction process should be natural. We need to ensure that data labels align with human perceptions of ’agentic’ and ’communal’ language.
@@
    The data construction process should be fair. Balancing measures should be taken to prevent potential biases from label imbalance. If the constructed data is built based on an original dataset, we also need to ensure there is no social bias propagation from the original data.
To collect a dataset with agentic and communal sentences through a mechanism that is scalable, consistent, natural, and fair, we adopt a novel dataset construction framework that consists of an automated component and a human-involved component. We begin by preprocessing a personal biography dataset (Lebret et al., 2016) into sentences, aiming at using these as seed texts to construct agentic and communal texts through paraphrasing. This step ensures the fairness of collected dataset, since (1) the raw data output would be balanced between the two labels, and (2) each sentence in each biography would have an agentic paraphrase and a communal paraphrase, preventing social bias propagation like having more agentic sentences for dominant social groups. Next, we adopt Openai’s gpt-3.5-turbo-1106 model (OpenAI, 2022) to paraphrase each sentence into an agentic version and a communal version. This ensures scalability through an automated generation pipeline, and also guarantees consistency since all paraphrases would come from a single source (in contrast with using human-written paraphrases, which is hard to scale and might result in drastically subjective writing tones). Furthermore, we utilize a human verification step to ensure the naturalness of the generated dataset. We invite $2$ human annotators, who are native speakers of English, to re-label each data and identify ambiguous cases. Finally, data entries with ambiguity are removed and ground truth labels of the LAC dataset are decided by a majority vote between the annotators’ labels and the paraphrasing target (i.e. whether a sentence was generated as an “agentic” or “communal” paraphrase). We elaborate on full details of dataset construction in Appendix B.
3.2.2 Dataset Statistics
@@ 3.2.3 Building A Language Agency Classifier With LAC
We experiment with both discriminative and generative models as base models for training language agency classifiers. Based on performances on LAC’s test set, we choose the fine-tuned BERT model as the language agency classifier in further experiments. Appendix C provides details of training and inferencing the classifiers, in which Table 11 reports classifier performances.
@@ 3.3 Quantifying Language Agency Bias in LLMs
We use the LAC-trained agency classifier to build quantitative metrics for measuring language agency bias in LLM generations. Specifically, we designed $2$ types of metrics: Intra-Group Agentic-Communal Ratio Gaps and Inter-Group Ratio Gap Variances. Intra-Group metrics objectively measure the agency level in texts generated for different demographic groups, whereas inter-group metrics estimate the variability of agency levels across groups.
@@ 4.1 Findings 1: LLM generations are More Gender Biased than Human-Written Texts
We establish comparison with bias in LLM-generated texts by incorporating analysis on 3 existing datasets: human-written biographies in Bias in Bios, human-written professor reviews on RateMyProfessor, and the reference letter dataset in  Wan et al. (2023)’s work, which consists of letters generated by LLMs given extensive biographical information (e.g. multi-sentence descriptions of career development) about specific individuals. Since we do not find any publicly available large-scale dataset for reference letters,  Wan et al. (2023)’s data is our best choice as a proxy of human-written letters. In addition, we did not find any openly-accessible datasets from the 3 categories that include racial information, limiting our analysis to gender biases.
Dataset Model Gender Diff. (M-F)
Biography Human $10.12$
ChatGPT $8.49$
Mistral 10.87
Llama3 $8.51$
Professor Review Human $1.86$
ChatGPT $6.57$
Mistral $8.14$
Llama3 $11.51$
Reference Letter Wan et al. (2023) $4.64$
ChatGPT $9.33$
Mistral $10.84$
Llama3 $9.44$
[…]
table Language agency gender bias in human-written and LLM-generated texts, measured by gender difference in agency-communal ratio gaps. Highest bias for each type of text is in bold.
4.1.1 Human-Written Texts: Dataset Details
We experiment with $3$ publicly accessible datasets of personal biographies, professor reviews, and reference letters. Full details of all datasets are in Appendix D.
Personal Biographies   We use Bias in Bios De-Arteaga et al. (2019), a biography dataset extracted from Wikipedia pages. Since the biography data for different professions are significantly imbalanced, we randomly sample $120$ biographies for each gender for each of the professions. A full list of professions in the pre-processed dataset is in Appendix D, Table 12.
Professor Reviews   We use an open-access sample dataset of student-written reviews for professors  [^1] https://github.com/x-zhe/RateMyProfessor_Sample_Dataset, which was web-crawled from the RateMyProfessor website  [^2] https://www.ratemyprofessors.com/. We first remove the majority of data entries without professors’ gender information. Since the remaining data is scarce and unevenly distributed across genders and departments, we remove data from departments with less than $10$ reviews for either gender. A full list of departments and corresponding gender distributions of professor reviews in the pre-processed dataset is provided in Appendix D, Table 13.
4.1.2 Comparison Results
Table 2 and Figure 2 show language agency gender biases in human-written and LLM-generated biographies, professor reviews, and reference letters. We report the gender differences (Male - Female) in the intra-group agency-communal ratio gaps. Below are our observations:
Biases observed in human-written texts in our study align with findings of social science studies. We stratify analysis on the human-written biography dataset based on professions in Appendix E, and found that occupations with greatest biases—such as pastor, architect, and software engineer—are also reported by real-world studies to be male-dominated (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Academic departments in which the highest language agency biases in professor reviews are identified—such as Accounting, Sociology, and Chemistry—have also been proven for male dominance (200, 2009; Girgus, ; Seijo, ). The alignment between our observations and real-world inequalities further validates the effectiveness of language agency in capturing social biases.
LLM-generated texts demonstrate more severe language agency gender biases than humans. As shown in Figure  2 and Table 2, for all 3 text categories, the highest gender bias levels, as measured by the gender differences in intra-group ratio gaps between agentic and communal sentences, are observed in LLMs. For professor reviews and reference letters, human-written texts demonstrate remarkably less bias than all 3 LLMs investigated. This warns of the potential propagation and even amplification of social biases in LLM-generated texts.
@@ 4.1.2 Comparison Results
Table 3: Experiment results for gender, racial, and intersectional bias in language agency of 3 investigated LLMs, across 3 text generation tasks. Highest bias level for each task for each LLM is underlined. Overall bias level across all tasks and all bias dimensions for each LLM is in bold. Llama3 demonstrates the highest overall bias in language agency (*).

### v4

@@ Abstract
Social biases can manifest in language agency. While several studies approached agency-related bias in human-written language, very limited research has investigated such biases in Large Language Model (LLM)-generated content. In addition, previous works often rely on string-matching techniques to identify agentic and communal words within texts, which fall short of accurately classifying language agency. We introduce the novel Language Agency Bias Evaluation (LABE) benchmark, which comprehensively evaluates biases in LLMs by analyzing agency levels attributed to different demographic groups in model generations. LABE leverages 5,400 template-based prompts, an accurate agency classifier, and corresponding bias metrics to test for gender, racial, and intersectional language agency biases in LLMs on 3 text generation tasks: biographies, professor reviews, and reference letters. We also contribute the Language Agency Classification (LAC) dataset, consisting of 3,724 agentic and communal sentences. Using LABE, we unveil language agency social biases in 3 recent LLMs: ChatGPT, Llama3, and Mistral. We observe that: (1) LLM generations tend to demonstrate greater gender bias than human-written texts; (2) Models demonstrate remarkably higher levels of intersectional bias than the other bias aspects. Those who are at the intersection of gender and racial minority groups—such as Black females—are consistently described by texts with lower levels of agency, aligning with real-world social inequalities; (3) Among the 3 LLMs investigated, Llama3 demonstrates the greatest overall bias; (4) Not only does prompt-based mitigation fail to resolve language agency bias in LLMs, but it frequently leads to the exacerbation of biases in generated texts.
White Men Lead, Black Women Help? Benchmarking Language Agency Social Biases in LLMs
Yixin Wan  and Kai-Wei Chang Department of Computer Science University of California, Los Angeles {elaine1wan, kwchang}@cs.ucla.edu
@@ 1 Introduction
Social biases manifest through the varying levels of agency in texts describing different demographic groups Grimm et al. (2020); Polanco-Santana et al. (2021); Stahl et al. (2022); Wan et al. (2023). For instance, bias exists in texts portraying demographic minority groups—such as Black individuals and women—as being communal (e.g. “warm” and “helpful”), and dominant social groups—such as White individuals and men—as being agentic (e.g. “authoritative” and“in charge of” things) Cugno (2020); Grimm et al. (2020). While a body of works in social science (Akos and Kretchmar, 2016; Grimm et al., 2020; Polanco-Santana et al., 2021; Park et al., 2021) and NLP (Sap et al., 2017; Ma et al., 2020; Park et al., 2021; Stahl et al., 2022; Wan et al., 2023) studied agency level in texts, they suffer from several remarkable drawbacks:
@@
2. Existing methods to measure language agency struggle with achieving accuracy and reliability. Prior works often utilized string matching for words in agentic and communal lexicons to measure agency. However, string matching and sentiment-based approaches only yield $46.65$ and $52.28$ in agency classification accuracy, respectively (as shown in Appendix B, Table 11). A qualitative example is provided in Figure 1: while differences in language agency are clearly observable in the texts, string matching yields 00 agentic and 00 communal words; sentiment classifier labels both texts as “positive”.  Wan et al. (2023) utilized a model-based method for measuring agency, but their model only achieves $66.49\%$ classification accuracy (Appendix B, Table 11).
@@ 2.1 Biases in Human-Written and LLM-generated Texts
The presence of gender, racial, and intersectional bias in human society has significantly impacted human language Blodgett et al. (2020); Doughman et al. (2021) and generative LLMs, which utilize extensive texts for training. We investigate biases in 3 different categories of texts: biographies, professor reviews, and reference letters.
Bias in BiographiesWagner et al. (2016); Field et al. (2022), and  Park et al. (2021) studied gender biases in Wikipedia biographies.  Park et al. (2021) analyzed biases in power, agency, and sentiment words in biography pages;  Wagner et al. (2016) revealed negative linguistic biases in womens’ pages.  Field et al. (2022) and  Adams et al. (2019) studied racial biases in editorial traits such as length and academic rank.  Field et al. (2022); Adams et al. (2019) and  Lemieux et al. (2023) stressed the importance of studying intersectional gender and racial biases in Wikipedia. Along similar lines, Otterbacher (2015) found biases towards Black female actresses in IMDB biographies.
Bias in Professor Reviews   Prior works (Roper, 2019; Macnell et al., 2014) have revealed gender biases in student ratings for professors—instructors with female perceived gender received lower ratings than males. Schmidt visualized the gendered language in RateMyProfessor reviews by string matching for gender-indicative words.  Reid (2010) showed that professors from racial minority groups received more negative RateMyProfessor evaluations.  Chávez and Mitchell (2020) further revealed intersectional biases towards female professors of racial minorities in professor reviews.
@@ 3 The Language Agency Bias Evaluation (LABE) Framework
In this paper, we propose the Language Agency Bias Evaluation (LABE) benchmark for comprehensively assessing language agency biases in LLMs across race, gender, and intersectional identities. LABE prompts LLMs to generate descriptive texts for multiple demographic groups, and assesses biases by inspecting the variability in language agency.
@@ 3.1 Generative Discriptive Texts for Demographic Groups with LLMs
Wan et al. (2023) proposed the Context-Less Generation (CLG) setting, in which they adopt templates and descriptors to prompt for a variety of LLM-generated reference letters for different genders. Inspired by CLG, we extend the setting to 3 different text generation tasks: biography, professor review, and reference letter generation. We combine descriptors with demographic information—such as race, gender, or intersectional identities—and template-based prompts to query for LLMs’ generation. Each prompt must contain race and gender descriptors. For the name descriptor, we prompt ChatGPT to generate 5 popular names for each gender and race intersectional group. Additional descriptors like occupation and department are included to improve prompt variability. The final LABE benchmark tests LLMs on 2,400 templated-based prompts for biography generation, 600 for professor review, and 2,400 for reference letters. Note that entry numbers differ due to the difference in descriptors used (departments for professor review, whereas occupations for the other 2). Full details are in Appendix A.
@@ 3.2 Evaluating Language Agency: The Language Agency Classification (LAC) Dataset
For building accurate automated evaluation tools for language agency, we propose the Language Agency Classification (LAC) dataset, a corpus with 3,724 agentic and communal sentences with corresponding labels. We adopt an efficient automated data generation pipeline and a verification step by English-speaking annotators.
@@ 3.2.1 Dataset Collection
    Scalablity. Since we are constructing a classification dataset, we need to ensure enough entries for agentic language and communal language to train a useful classifier.
@@
    Consistency. Quality of texts for all entries should remain consistent, and should not differ from part to part.
@@
    Naturalness. We need to ensure that data labels align with human perceptions of “agentic” and “communal” language.
@@
    Fairness. Balancing measures should be taken to prevent potential biases from label imbalance or bias propagation from a seed dataset.
To collect a dataset with agentic and communal sentences through a mechanism that is scalable, consistent, natural, and fair, we adopt a novel dataset construction framework that consists of an automated component and a human-involved component. We begin by preprocessing a personal biography dataset (Lebret et al., 2016) into sentences, aiming at using these as seed texts to construct agentic and communal texts through paraphrasing. This step ensures the fairness of collected dataset, since (1) the raw data output would be balanced between the two labels, and (2) each sentence in each biography would have an agentic paraphrase and a communal paraphrase, preventing social bias propagation like having more agentic sentences for dominant social groups. Next, we adopt Openai’s gpt-3.5-turbo-1106 model (OpenAI, 2022) to paraphrase each sentence into an agentic version and a communal version. This ensures scalability through an automated generation pipeline, and also guarantees consistency since all paraphrases would come from a single source (in contrast with using human-written paraphrases, which is hard to scale and might result in drastically subjective writing tones). Furthermore, we utilize a human verification step to ensure the naturalness of the generated dataset. We invite $2$ human annotators, who are native speakers of English, to re-label each data and identify ambiguous cases. Finally, data entries with ambiguity are removed and ground truth labels of the LAC dataset are decided by a majority vote between the annotators’ labels and the paraphrasing target (i.e. whether a sentence was generated as an “agentic” or “communal” paraphrase). Full details of dataset construction are in Appendix B. Details on dataset statistics are in Appendix B.5
@@ 3.2.2 Building A Language Agency Classifier With LAC
We experiment with both discriminative and generative models as base models for training language agency classifiers. Based on performances on LAC’s test set, we choose the fine-tuned BERT model as the language agency classifier in further experiments. Appendix B provides details of training and inferencing the classifiers, in which Table 11 reports classifier performances.
@@ 3.3 Quantifying Language Agency Bias in LLMs
We use the LAC-trained agency classifier to build quantitative metrics for measuring language agency bias in LLM generations. Specifically, we compute the Intra-Group Agentic-Communal Ratio Gaps as the objective agency level, and measure biases through Inter-Group Ratio Gap Variances. We establish the inter-group variance as our bias evaluation metric, since it assesses the variability of agency levels across social groups.
@@ 4 Unveiling Language Agency Biases in LLMs with LABE
Table 1: Experiment results for gender, racial, and intersectional bias in language agency of 3 investigated LLMs, across 3 text generation tasks. Greatest bias for each task for each LLM is underlined. Overall bias level across all tasks and all bias dimensions for each LLM is in bold. Llama3 demonstrates the highest overall agency bias (*).
We establish comparison with bias in LLM-generated texts by incorporating analysis on 3 existing datasets: human-written biographies in Bias in Bios, human-written professor reviews on RateMyProfessor, and the reference letter dataset in  Wan et al. (2023)’s work, which consists of letters generated by LLMs given extensive biographical information (e.g. multi-sentence descriptions of career development) about specific individuals. Since we do not find any publicly available large-scale dataset for reference letters,  Wan et al. (2023)’s data is our best choice as a proxy of human-written letters. Additionally, no openly-accessible datasets with racial information were found in our search, limiting our analysis to gender biases.
4.1.1 Human-Written Texts: Dataset Details
We experiment with $3$ publicly accessible datasets of personal biographies, professor reviews, and reference letters. Full details of all datasets are in Appendix C.
Personal Biographies   We use Bias in Bios De-Arteaga et al. (2019), a biography dataset extracted from Wikipedia pages. Since the biography data for different professions are significantly imbalanced, we randomly sample $120$ biographies for each gender for each of the professions. A full list of professions in the pre-processed dataset is in Appendix C, Table 12.
Professor Reviews   We use an open-access sample dataset of student-written reviews for professors  [^1] https://data.mendeley.com/datasets/fvtfjyvw7d/2, which was web-crawled from the RateMyProfessor website  [^2] https://www.ratemyprofessors.com/. We first remove the majority of data entries without professors’ gender information. Since the remaining data is scarce and unevenly distributed across genders and departments, we remove data from departments with less than $10$ reviews for either gender. A full list of departments and corresponding gender distributions of professor reviews in the pre-processed dataset is provided in Appendix C, Table 13.
4.1.2 Comparison Results
Figure 2 visualizes language agency gender biases in human-written and LLM-generated biographies, professor reviews, and reference letters. We report the gender differences (Male - Female) in the intra-group agency-communal ratio gaps. Quantitative results are in Appendix D, Table 14. Below are our observations:
Biases observed in human-written texts in our study align with findings of social science studies. We stratify analysis on the human-written biography dataset based on professions in Appendix D, and found that occupations with greatest biases—such as pastor, architect, and software engineer—are also reported by real-world studies to be male-dominated (Kathleen Schubring, ; A. Nicholson et al., ; Kaminski, ). Academic departments in which the highest language agency biases in professor reviews are identified—such as Accounting, Sociology, and Chemistry—have also been proven for male dominance (200, 2009; Girgus, ; Seijo, ). Alignment between our observations and real-world inequalities further shows the effectiveness of agency in capturing social biases.
LLM-generated texts demonstrate more severe language agency gender biases than humans. As shown in Figure  2, for all 3 text categories, the highest gender bias levels, as measured by the gender differences in intra-group ratio gaps between agentic and communal sentences, are observed in LLMs. For professor reviews and reference letters, human-written texts demonstrate remarkably less bias than LLMs. This warns of the potential propagation and even amplification of social biases in LLM-generated texts.